OpenStudy (doc.brown):

What's the kinetic energy of a 12.0g particle with a velocity of 400m/s?

OpenStudy (anonymous):

0.5 x mass x vel^2= 0.5 x .12 x400 x 400 joules

OpenStudy (doc.brown):

Is this right?\[KE=\frac{1}{2}mv^2\]\[KE=\frac{1}{2}(12.0g)(400\frac{m}{s})^2\]\[9.6\times10^5\frac{gm^2}{s^2}\]\[9.6\times10^2\frac{kgm^2}{s^2}\]\[9.6\times10^2J\]

OpenStudy (doc.brown):

Thanks mr.singh for the formula, however, \[12g\times\frac{kg}{1000g}=0.012kg\]not 0.12

OpenStudy (loser66):

is it not solved?

OpenStudy (doc.brown):

I was just waiting for a "yes this looks right to me"

OpenStudy (loser66):

ok, so, "yes, this looks good to me" lol